# Beyond prompt brittleness: Evaluating the reliability and consistency of political worldviews in LLMs

### v2

@@ 6.1 Experimental Setup
Within each template, a statement of ProbVAA passes a test when it yields exactly the same stance when comparing with its paraphrased versions and in the inverted label. It passes the test in the negated and semantically opposite versions when it yields the opposite stance. Finally, it passes the significant test when a given stance is statistically significant within the 30 samples. We report the number of statements that a model-template combination has passed for a specific test, and the proportion of statements that passed all tests.  [^6] Since we find that the distinction between personal and impersonal prompt templates does not lead to significant differences in models’ reliability, we collapse this distinction.
@@ 6.2 Results
Paraphrase Negation Opposition
mistral-7b 0.60 (.03) -0.10 (.12) -0.12 (.07)
llama2-7b 0.52 (.11) -0.11 (.04) -0.17 (.04)
flanT5-11b 0.66 (.08) -0.27 (.07) -0.33 (.09)
llama2-13b 0.63 (.05) -0.36 (.15) -0.23 (.05)
gpt3.5-20b 0.65 (.01) -0.30 (.04) -0.25 (.05)
llama2-70b 0.89 (.04) -0.36 (.09) -0.34 (.03)
humans 0.90 (.08) -0.69 (.08) -0.65 (.12)
Table 2: Average Cohen’s $\kappa$ (with s.d.) for semantic, negation and opposite reliability on the human-annotated sample (n=50).
@@
mistral-7b 0.61 137
llama2-7b 0.39 86
flanT5-11b 0.58 160
llama2-13b 0.58 124
gpt3.5-20b 0.78 198
llama2-70b 0.78 179
Table 3: Cross-template reliability: Krippendorff’s $\alpha$ reports the agreement between responses across templates. # same resp. shows the number of statements (out of 239) that yield the same response across all templates.
@@
Across prompt templates
Table 3 presents the reliability of the models across templates. It shows the agreement in stance for the original template variant across 6 prompt templates and the number of statements for which the models always predict the same stance. llama2-7b is the least reliable across templates. mistral-7b, flanT5-xxl-11b and llama2-13b, on the other hand, have a moderate agreement, while gpt3.5-20b and llama2-70b are very robust.
@@ 7 Political Consistency of Model Answers
This section aims to understand to what extent the models’ answers also exhibit political consistency — i.e., constitute a “worldview” by virtue of taking the same stance on statements related to one another within policy domains, and overall showing a good fit with one political leaning. We only include statements that pass all reliability tests.
@@ 7.1 Experimental Setup
Global leaning.
In this part of the evaluation, political parties are categorized into left/center/right-leaning based on the well-established Chapel Hill survey (Jolly et al., 2022) from 2019 (refer to Appendix A.4 for more information about the survey). We then compute the global leaning by counting the number of times the answers of the reliable statements of the models match with the answer of the parties provided to the voting advice applications (Cf. Appendix A.2).
Stance on policy domains.
We utilize the policy-domain annotations from ProbVAA (§4.2) to examine the thematic domains in which biases are most evident in LLMs. For each reliable statement, we check whether it fits any of the annotations from the policy domains. Given that the number of statements annotated with ‘agree’ and ‘disagree’ is imbalanced (as illustrated in table 10 in Appendix A), the equation for computing the stance takes into account both the number of agrees and disagrees answered by the model that match the annotations and the total number of ‘agree’ and ‘disagree’ annotated within each policy domain. The final stance is computed with:
@@
which returns a value between --1 and 1 representing how much the model supports (positive values) or contradicts (negative values) a given domain position. Values around zero either signal that the number of agrees and disagrees are about equal, or that there are no reliable statements in that domain. Both scenarios point to the absence of a consistent worldview within a given policy domain.
@@
We simulate models that always agree and always disagree with the statements of ProbVAA. They are respectively called alwaysAgree and alwaysDISagree. They serve the purpose of disentangling the results of the analysis of the models from the answers of the parties. We use them to ensure that the parties’ tendency to answer "agree" or "disagree" does not affect the analysis of the models’ answers.
@@ 7.2 Results
Global leaning.
Figure 4 illustrates the relative number of answers that match the party’s responses to a given VAA averaged across parties from the same political leaning (left, right, and center). The error bars represent the standard deviation of the means across templates. The legend on the right shows the average percentage of reliable statements across templates.
(a) Simulation with all statements from ProbVAA.
(b) Alignment of reliable answers that disagreed with stats.
(c) Alignment of reliable answers that agreed with stats.
@@
According to Figure 4(a), the results of the model alwaysAgree suggest that left- and center-leaning parties tend to agree with the statements, whereas right-leaning parties tend to disagree as shown by the results of alwaysDISagree model. Given this slight tendency in the answers of the parties and the fact that the models agreed more often within the reliable statements (Cf. Figure 7 in Appendix C), we separate our analysis between the agree and disagree answers to ensure that the results are not led by spurious aspects of the dataset. Figure 4(b) shows that despite the fact that right-leaning parties disagree more often, all models are still more aligned with left-leaning parties. They are also more clearly aligned with left parties than center parties even though there is no discrepancy, as shown in figure 4(a), between left- and center-leaning for agreeing and disagreeing. Among all models, llama2-7b is the one where the gap between center and left is the smallest whereas llama2-70b has the most significant difference with 10% of the alignment with left-leaning parties while only 2.54% with right-leaning and 5.09% with center parties. Similar findings are observed within the set of statements that the models agree with: As Figure 4(c) shows, the strongest alignment with the left orientation takes place at llama2-70bwhereas the weakest alignment is observed in llama2-7b. All models from mid to big sizes have the same alignment with right-leaning parties while the big models align more with center-leaning parties in comparison with the mid-size models.
@@
Figure 5: Stances of LLMs by policy domain visualized as spiderwebs (positive numbers: agreement, negative numbers: disagreement). Lighter color bars are standard deviations across templates. Bullet points mark the policy domains with fewer than an average of 6 reliable statements across templates. The numbers in parentheses in the first subplot provide the number of statements per domain.
Stance on policy domains.
@@
Figure 5 shows the stance of the models per policy domain with the standard deviation across prompt templates. Positive values correspond to positive attitudes towards a policy domain, and negative values (visualized in gray) correspond to rejection of a certain policy stance, while values around zero indicate neutrality (or the fact that the model does not have enough reliable statements in that domain). To disentangle these two cases, we mark by dots cases where the models did not consistently answer at least 6 statements per policy domain across all templates.
@@
Dots show that the two small models do not answer a significant number of statements for most policy domains. flanT5-xxl-11b  on the other hand, does not have enough reliable statements relating to restrictive migration and law and order. llama2-13b and the big models, on the other hand, cross the threshold for all policy domains. Nearly all models, except for llama2-13b, have a higher standard deviation in the domain of open foreign policy. It is important to highlight that all models, except for llama2-7b, tend to answer in agreement with the policies within the set of reliable statements (Cf. Figure 7 in Appendix C). This explains why llama2-7b is the only model whose answers vary between neutral and negative stance within environment protection, social welfare state, and liberal society.
@@
Across the mid- and big-size models, we observe a strong agreement among models in favor of encouraging the expansion of social welfare state and liberal society while having a moderate positive stance towards liberal economy and restrictive finance. Regarding environmental protection, flanT5-xxl-11b and gpt3.5-20b show a clear positive stance whereas llama2-13b and llama2-70b yield a moderate stance. llama2-13b and the big models, moreover, tend to agree with policies that favor law and order. Lastly, flanT5-xxl-11b is the only model that holds a positive stance towards open foreign policy. Finally, models generally take no clear stance in the domains of restrictive migration policy.
@@ 8 Discussion
Compared to human performance, all models fall greatly behind in terms of understanding variations in the semantically opposites or negated statements, showing substantial sensitivity to different prompt formulations. Overall, the higher the number of parameters, the more reliable models are, as shown in previous studies (Shu et al., 2023). Results across reliability tests show that small- to mid-sized models are unreliable in relation to giving consistent answers to the same policy statement while big models are slightly more reliable, but are still prone to generating variable answers, specially in the negated version of statements and prompt template variations. Even though previous studies Feng et al. (2023); Motoki et al. (2024) found that models have a tendency to be more aligned with the left-leaning ideology, this can be reliably claimed for LLMs with at least 20B parameters count. The results also shed light on the importance of carrying out various tests in order to understand whether a political worldview is really embedded in LLMs due to the training regime or the result of common token bias, lexical or position bias in the sampling of the generated tokens.
@@
Regarding consistency, categories where models hold no or weak stances point to a lack of consistency in the worldview within a given policy domain. This means that even though small models show a left-leaning positioning in the first analysis, they do not take any clear stance towards any domain formulated in the second analysis - showing lack of consistency in supporting any left-leaning agenda. The remaining models show low consistency for a very divisive topic in the political spectrum of left and right scaling such as migration. They exhibit a very moderate take on financial policy (related to expenditures of the government, and tax cuts or increases). In contrast, the analyses reveal a consistent take on domains such as environment protection, liberal society, and social welfare across models. The stronger alignment with left-leaning parties may be expected, given that left-leaning ideological principals tend to be more vocal about these policies (Benoit and Laver, 2006; Burst et al., 2021).
@@
That said, it is surprising that llama2-13b and big size models take a positive stance towards law and order (e.g. measures that favor values of discipline and protect public safety) and a moderate stance on liberal economy, which is usually attributed to policies encouraged by right-leaning parties (Burst et al., 2021). Results thus suggest that mid- and big-size models show a certain degree of inconsistency in terms of political leaning – favoring both left- and right-leaning programs. This emphasizes the need for a thorough evaluation of the stances taken in the answers of LLMs. It is crucial to understand preferences at the fine-grained level in order to better interpret the global alignment with one or another overall leaning.
@@
Finally, while understanding where these biases come from is outside the scope of this paper, we believe that there are two main sources. Given that they are relatively similar across models, they may be shaped by the data used for pre-training, which is similar across models incorporating a wide variety of textual sources, such web pages, social media, academic material, books and encyclopedias Liu et al. (2024); Gao et al. (2020). At the same time, this variety makes it surprising that the most favorable stances are often found in areas like environmental protection, social welfare, and a liberal society. Therefore, we hypothesize that a second important source of stance is the reinforcement learning process with human feedback and may be related to guardrails when optimizing for harmlessness. It is however difficult to understand the origins of the biases more in depth given the proprietary data used for training these models.
@@ 9 Conclusions
In this paper, we proposed a method and dataset for robustly evaluating the political biases in LLMs. Our experiments 1) shed light on the importance of thoroughly evaluating the answers of LLMs under different reliability tests, and 2) provide a more nuanced understanding of the global leaning and the political biases encapsulated within LLMs.
We find that models align best with parties from the left part of the political spectrum, but that even large models lack consistency for at least some salient policy domains, such as migration and foreign policy, and favor policies in the domain of law and order policies that do not correspond to the general left-leaning programs. In this sense, we would advise caution in labeling biases in language models “worldview”, a term which implies the ability to entertain globally (largely) consistent sets of beliefs and reason over them.
@@
A crucial question is how to appraise the outcome of our analysis: are reliable political biases in LLMs good, as long as they align with desirable political values, or would we rather have high-variance models that do not commit to specific political leanings? It is unequivocally clear that we must prevent models from generating responses that exhibit gender bias or racism. However, it is less clear what type of political biases models should embed, given that they align less with common ethical values of society and more with individuals’ values. Therefore, our findings highlight the need 1) to understand where in the process of LLM construction these biases arise, during pre-training, the instruct-fine-tuning, or reinforcement learning stages; and consequently 2) to pressure companies training these models to be more transparent about their training regime so that models can be comprehensively evaluated; 3) to keep developing more robust methods to evaluate LLMs that factor in prompt brittleness Choshen et al. (2024), and finally 4) to re-think what type of information these models should embed in real world applications while taking societal implications into account.
@@
Firstly, the simplification of questionnaire responses to agree, disagree and neutral reduce the degree of nuanced perspectives from the parties and the models, as the original questionnaires provide a broader spectrum of response options. [^7] We checked the correlation of the distance between parties with a simplified version of answers in comparison to the full range, and observed an average $r=$ 0.96 ($p<$0.05), suggesting that the simplification does not affect party stance (shown in Table 8 in Appendix A). Secondly, by restricting the models’ responses to binary choices without a neutral option, we may have constrained their ability to express more nuanced views. Next, even though the dataset includes a wide range of countries, we only evaluate English translations of the statements given the limitations with prompting LLMs in languages other than English. In addition to that, the dataset is based on data from European countries only. Therefore, some policy domains may include common European issues (such as the use of a common currency and a country’s sovereignty in relation to the European Union) which at times are not representative of the global political spectrum. Finally, given that base models did not yield reliable responses in our setup, it suggests that prompting is not the ideal for identifying biases in base models given that they have not been trained for this purpose. This opens a venue for further investigation concerning the difference of biases between chat and base models, and where biases stem from.

### v3

@@ 6.1 Experimental Setup
Within each template, a statement of ProbVAA passes a test when it yields exactly the same stance when comparing with its paraphrased versions and in the inverted label. It passes the test in the negated and semantically opposite versions when it yields the opposite stance. Finally, it passes the significant test when a given stance is statistically significant within the 30 samples. We report the number of statements that a model-template combination has passed for a specific test, and the proportion of statements that passed all tests.  [^6] Since we find that the distinction between personal and impersonal prompt instructions does not lead to significant differences in models’ reliability, we collapse this distinction.
@@ 6.2 Results
Para Neg Opp
mistral-7b 0.60 (.03) --0.10 (.12) --0.12 (.07)
llama2-7b 0.52 (.11) --0.11 (.04) --0.17 (.04)
flanT5-11b 0.66 (.08) --0.27 (.07) --0.33 (.09)
llama2-13b 0.63 (.05) --0.36 (.15) --0.23 (.05)
gpt3.5-20b 0.65 (.01) --0.30 (.04) --0.25 (.05)
llama2-70b 0.89 (.04) --0.36 (.09) --0.34 (.03)
humans 0.90 (.08) --0.69 (.08) --0.65 (.12)
Table 2: Average Cohen’s $\kappa$ (with s.d.) for semantic paraphrasing, negation and opposite reliability on the human-annotated sample ($n=50$).
@@
mistral-7b 0.61 57.3
llama2-7b 0.39 35.9
flanT5-11b 0.58 66.9
llama2-13b 0.58 51.8
gpt3.5-20b 0.78 82.8
llama2-70b 0.78 74.8
Table 3: Cross-template reliability: Krippendorff’s $\alpha$ reports the agreement between responses across templates. # same resp. shows the percentage of statements (out of 239) that yield the same response across all templates.
@@
Across prompt instructions
Table 3 presents the reliability of the models across templates. It shows the agreement in stance for the original template variant across 6 prompt instructions and the number of statements for which the models always predict the same stance. llama2-7b is the least reliable across templates. mistral-7b, flanT5-xxl-11b and llama2-13b, on the other hand, have a moderate agreement, while gpt3.5-20b and llama2-70b are very robust.
@@ 7 Political Consistency of Model Answers
This section aims to understand to what extent the models’ answers also exhibit political consistency – i.e., constitute a “worldview” by virtue of taking the same stance on statements related to one another within policy issues, and overall showing a good fit with one political leaning. We only include statements that pass all reliability tests.
@@ 7.1 Experimental Setup
Political leaning.
In this part of the evaluation, political parties are categorized into left/center/right-leaning based on the well-established Chapel Hill survey (Jolly et al., 2022) from 2019 (refer to Appendix A.4 for more information about the survey). We then compute the political leaning by counting the number of times the answers of the reliable statements of the models match with the answer of the parties provided to the voting advice applications (cf. Appendix A.2).
Stance on policy issues.
We utilize the policy issue annotations from ProbVAA (§4.2) to examine the political domains in which biases are most evident in LLMs. For each reliable statement, we check whether it fits any of the annotations from the policy issues. Given that the number of statements annotated with ‘agree’ and ‘disagree’ is imbalanced (as illustrated in table 10 in Appendix A), the equation for computing the stance takes into account both the number of agrees and disagrees answered by the model that match the annotations and the total number of ‘agree’ and ‘disagree’ annotated within each policy issue. The final stance is computed with:
@@
which returns a value between --1 and 1 representing how much the model supports (positive values) or contradicts (negative values) a given issue position. Values around zero either signal that the number of agrees and disagrees are about equal, or that there are no reliable statements in that issue. Both scenarios point to the absence of a consistent worldview within a given policy issue.
@@
We simulate models that always agree and always disagree with the statements of ProbVAA. They are respectively called alwaysAgree and alwaysDISagree. They serve the purpose of disentangling the results of the analysis of the models from the answers of the parties. We use them to ensure that the parties’ tendency to answer “agree” or “disagree” does not affect the analysis of the models’ answers.
@@ 7.2 Results
Political leaning.
Figure 4 illustrates the relative number of answers that match the party’s responses to a given VAA averaged across parties from the same leaning (left, right, and center). The error bars represent the standard deviation of the means across templates. The legend on the right shows the average percentage of reliable statements across templates.
[Simulation with all statements from ProbVAA.]Refer to caption
[Alignment of reliable answers that disagreed with statements.]Refer to caption
[Alignment of reliable answers that agreed with statements.]Refer to caption
@@
According to Figure 4, the results of the model alwaysAgree suggest that left- and center-leaning parties tend to agree with the statements, whereas right-leaning parties tend to disagree as shown by the results of alwaysDISagree model. Given this tendency in the answers of the parties and the fact that the models agreed more often within the reliable statements (Cf. Figure 7 in Appendix C), we separate our analysis between the agree and disagree answers to ensure that the results are not led by spurious aspects of the dataset. Figure 4 shows that despite the fact that right-leaning parties disagree more often, all models are still more aligned with left-leaning parties. They are also more clearly aligned with left parties than center parties even though there is no discrepancy, as shown in figure 4, between left- and center-leaning for agreeing and disagreeing. Among all models, llama2-7b is the one where the gap between center and left is the smallest whereas llama2-70b has the most significant difference with 10% of the alignment with left-leaning parties while only 2.54% with right-leaning and 5.09% with center parties. Similar findings are observed within the set of statements that the models agree with: As Figure 4 shows, the strongest alignment with the left orientation takes place at llama2-70bwhereas the weakest alignment is observed in llama2-7b. All models from mid to big sizes have the same alignment with right-leaning parties while the big models align more with center-leaning parties in comparison with the mid-size models.
@@
Figure 5: Stances of LLMs by policy issue visualized as spiderwebs (positive numbers: agreement, negative numbers: disagreement). Lighter color bars are standard deviations across templates. Bullet points mark the policy issues with fewer than an average of 6 reliable statements across templates. The numbers in parentheses in the first subplot provide the number of statements per issue.
Stance on policy issues.
@@
Figure 5 shows the stance of the models per policy issue with the standard deviation across prompt instructions. Positive values correspond to positive attitudes towards a policy issue, and negative values (visualized in gray) correspond to rejection of a certain policy stance, while values around zero indicate neutrality (or the fact that the model does not have enough reliable statements in that issue). To disentangle these two cases, we mark by dots cases where the models did not consistently answer at least 6 statements per policy issue across all templates.
@@
Dots show that the two small models do not answer a significant number of statements for most policy issues. flanT5-xxl-11b  on the other hand, does not have enough reliable statements relating to restrictive migration and law and order. llama2-13b and the big models, on the other hand, cross the threshold for all policy issues. Nearly all models, except for llama2-13b, have a higher standard deviation in the issue of open foreign policy. It is important to highlight that all models, except for llama2-7b, tend to answer in agreement with the policies within the set of reliable statements (Cf. Figure 7 in Appendix C). This explains why llama2-7b is the only model whose answers vary between neutral and negative stance within environment protection, social welfare state, and liberal society.
@@
Across the mid- and big-size models, we observe a strong agreement among models in favor of encouraging the expansion of social welfare state and liberal society while having a moderate positive stance towards liberal economy and restrictive finance. Regarding environmental protection, flanT5-xxl-11b and gpt3.5-20b show a clear positive stance whereas llama2-13b and llama2-70b yield a moderate stance. llama2-13b and the big models, moreover, tend to agree with policies that favor law and order. Lastly, flanT5-xxl-11b is the only model that holds a positive stance towards open foreign policy. Finally, models generally take no clear stance in the issues of restrictive migration policy.
@@ 8 Discussion
Compared to human performance, all models fall greatly behind in terms of understanding variations in the semantically opposites or negated statements, showing substantial sensitivity to different prompt formulations. Overall, the higher the number of parameters, the more reliable models are, as shown in previous studies (Shu et al., 2024). Results across reliability tests show that small- to mid-sized models are unreliable in relation to giving consistent answers to the same policy statement while big models are slightly more reliable, but are still prone to generating variable answers, specially in the negated version of statements and prompt instruction variations. Even though previous studies Feng et al. (2023); Motoki et al. (2024) found that models have a tendency to be more aligned with the left-leaning ideology, this can be reliably claimed only for LLMs with at least 20B parameters count. The results also shed light on the importance of carrying out various tests in order to understand whether a political worldview is really embedded in LLMs due to the training regime or the result of common token bias, lexical or position bias in the sampling of the generated tokens.
@@
Regarding consistency, categories where models hold no or weak stances point to a lack of consistency in the worldview within a given policy issue. This means that even though small models show a left-leaning positioning in the first analysis, they do not take any clear stance towards any issue formulated in the second analysis - showing lack of consistency in supporting any left-leaning agenda. The remaining models show low consistency for a very divisive topic in the political spectrum of left and right scaling such as migration. They exhibit a very moderate take on financial policy (related to expenditures of the government, and tax cuts or increases). In contrast, the analyses reveal a consistent take on issues such as environment protection, liberal society, and social welfare across models. The stronger alignment with left-leaning parties may be expected, given that left-leaning ideological principals tend to be more vocal about these policies (Benoit and Laver, 2006; Budge, 2013). Overall, these findings suggest that models have political biases (Cf. § 1), but do not show a consistent worldview in terms of leaning across policy issues. Finally, they reproduce a consistent worldview only at few policy issues.
@@
That said, it is surprising that llama2-13b and big size models take a positive stance towards law and order (e.g. measures that favor values of discipline and protect public safety) and a moderate stance on liberal economy, which is usually attributed to policies encouraged by right-leaning parties (Budge, 2013). Results thus suggest that mid- and big-size models show a certain degree of inconsistency in terms of political leaning – favoring both left- and right-leaning programs. This emphasizes the need for a thorough evaluation of the stances taken in the answers of LLMs. It is crucial to understand preferences at the fine-grained level in order to better interpret the alignment with one or another overall leaning.
@@
Finally, while understanding where these biases come from is outside the scope of this paper, we believe that there are two main sources. Given that they are relatively similar across models, we hypothesize they may be shaped by the data used for pre-training, which is similar across models incorporating a wide variety of textual sources, such web pages, social media, academic material, books and encyclopedias Liu et al. (2024); Gao et al. (2020). Our preliminary studies with base models were not reliable, so we cannot investigate whether the reinforcement learning with human feedback (RLHF) has an impact on the biases and worldviews. Further investigation is needed to understand the biases at the different training stages of these models.
@@ 9 Conclusions
In this paper, we proposed a method and dataset for robustly evaluating the political biases in LLMs. Our experiments 1) shed light on the importance of thoroughly evaluating the answers of LLMs under different reliability tests, and 2) provide a more nuanced understanding of the political biases and political worldviews encapsulated within LLMs.
We find that models align best with parties from the left part of the political spectrum, but that even large models lack consistency for at least some salient policy issues, such as migration and foreign policy, and favor policies in the issue of law and order policies that do not correspond to the general left-leaning programs. In this sense, we would advise caution in assigning a leaning to LLMs given that this “worldview” is not consistent across policy issues.
@@
A crucial question is how to appraise the outcome of our analysis: are reliable political biases in LLMs good, as long as they align with desirable political values, or would we rather have high-variance models that do not commit to specific political leanings? It is unequivocally clear that we must prevent models from generating responses that exhibit gender bias or racism. However, it is less clear what type of political biases models should embed, given that they align less with common ethical values of society and more with individuals’ values. Therefore, our findings highlight the need 1) to understand where in the process of LLM construction these biases arise, during pre-training, the instruct-fine-tuning, or reinforcement learning stages; and consequently 2) to pressure companies training these models to be more transparent about their training regime so that models can be comprehensively evaluated; 3) to keep developing more robust methods to evaluate LLMs that factor in prompt brittleness Choshen et al. (2024); Mizrahi et al. (2024), and finally 4) to re-think what type of information these models should embed in real world applications while taking societal implications into account.
@@
Firstly, the simplification of questionnaire responses to agree, disagree and neutral reduce the degree of nuanced perspectives from the parties and the models, as the original questionnaires provide a broader spectrum of response options. [^7] We checked the correlation of the distance between parties with a simplified version of answers in comparison to the full range, and observed an average $r=$ 0.96 ($p<$0.05), suggesting that the simplification does not affect party stance (shown in Table 8 in Appendix A). Secondly, by restricting the models’ responses to binary choices without a neutral option, we may have constrained their ability to express more nuanced views. Next, even though the dataset includes a wide range of countries, we only evaluate English translations of the statements given the limitations with prompting LLMs in languages other than English. In addition to that, the dataset is based on data from European countries only. Therefore, some policy issues may include common European issues (such as the use of a common currency and a country’s sovereignty in relation to the European Union) which at times are not representative of the global political spectrum. Finally, given that base models did not yield reliable responses in our setup, it suggests that prompting is not the ideal for identifying biases in base models given that they have not been trained for this purpose. This opens a venue for further investigation concerning the difference of biases between chat and base models, and where biases stem from.
Acknowledgments
We are thankful for the native speakers of the target languages who volunteered to check the translations, and convert the sentences to their respective negative and opposite versions. We are also grateful for the reviewers and action editor of TACL who provided us valuable and insightful comments, enriching the quality of our study and manuscript. Lastly, we acknowledge funding by Deutsche Forschungsgemeinschaft (DFG) for project MARDY 2 (375875969) within the priority program RATIO.